# Reinforcement Learning with Hidden Markov Models for Discovering Decision-Making Dynamics

### v1

@@ Abstract
Major depressive disorder (MDD), a leading cause of years of life lived with disability, presents challenges in diagnosis and treatment due to its complex and heterogeneous nature. Emerging evidence indicates that reward processing abnormalities may serve as a behavioral marker for MDD. To measure reward processing, patients perform computer-based behavioral tasks that involve making choices or responding to stimulants that are associated with different outcomes, such as gains or losses in the laboratory. Reinforcement learning (RL) models are fitted to extract parameters that measure various aspects of reward processing (e.g., reward sensitivity) to characterize how patients make decisions in behavioral tasks. Recent findings suggest the inadequacy of characterizing reward learning solely based on a single RL model; instead, there may be a switching of decision-making processes between multiple strategies. An important scientific question is how the dynamics of learning strategies in decision-making affect the reward learning ability of individuals with MDD. Motivated by the probabilistic reward task (PRT) within the EMBARC study, we propose a novel RL-HMM framework for analyzing reward-based decision-making. Our model accommodates learning strategy switching between two distinct approaches under a hidden Markov model (HMM): subjects making decisions based on the RL model or opting for random choices. We account for continuous RL state space and allow time-varying transition probabilities in the HMM. We introduce a computationally efficient EM algorithm for parameter estimation and employ a nonparametric bootstrap for inference. Extensive simulation studies validate the finite-sample performance of our method. We apply our approach to the EMBARC study to show that MDD patients are less engaged in RL compared to the healthy controls, and engagement is associated with brain activities in the negative affect circuitry during an emotional conflict task. reinforcement learning; reward tasks; state-switching; behavioral phenotyping; mental health; brain-behavior association
@@ 1 Introduction
Despite extensive research and clinical efforts dedicated to developing pharmacological and behavioral treatments for mental disorders over the years, a significant setback is symptom-based diagnoses that assume discrete disease categories, which may contribute to low remission rates and inadequate treatment responses as it does not fully account for the heterogeneity and complexity of mental disorders (Rush et al., 2006). In response to the limitations inherent in traditional symptom-based diagnoses and to address significant between-patient heterogeneity, the National Institute of Mental Health (NIMH) advocates for a paradigm shift through the Research Domain Criteria (RDoC) initiative (Insel et al., 2010). This shift involves redefining mental disorders by identifying latent constructs derived from biological and behavioral measures across diverse domains of functioning (e.g., positive/negative affect, decision-making, social processing) at different levels (e.g., cells, genes, and brain circuits), instead of solely relying on the clinical symptoms. Behavioral data collected from tasks such as the probabilistic reward task (PRT; Pizzagalli et al., 2005) and the emotion conflict task (Etkin et al., 2006) is a highly valued resource within the framework of RDoC. Utilizing carefully designed behavioral tasks with appropriate data- or theory-driven computational methods (Huys et al., 2016), researchers can extract valuable behavioral markers that serve as indicators of mental health conditions. Moreover, there is a critical need to examine the associations between the brain and behavior, as demonstrated by Fonzo et al. (2019). This exploration facilitates understanding the underlying brain mechanisms during behavioral tasks, allowing for the identification of brain abnormalities related to mental disorders, and ultimately contributes to refining diagnostic frameworks and developing more effective therapeutic interventions.
@@
Reward learning and decision-making are pivotal in the etiology of mental function (Kendler et al., 1999). Reward processing refers to how individuals perceive, value, and pursue rewards (e.g., money, social approval). MDD patients may show abnormalities in reward processing, such as reduced motivation, anhedonia (loss of pleasure), or reduced sensitivity to positive rewards. To measure reward processing, patients perform computer-based behavioral tasks that involve making choices or responding to stimulants that are associated with different outcomes, such as gains or losses in the laboratory. We focus on characterizing human decision-making behaviors through reward tasks (e.g., PRT), wherein individuals make decisions by interacting with a reward-based system and continually adapting their decision-making mechanisms based on obtained rewards.
@@
Biological evidence suggests that fluctuations in human dopaminergic neurons indicate errors in predicting future rewarding events (Schultz et al., 1997). This supports the use of simple prediction error learning (comparing real and expected reward) to update the value functions of different choices (Rescorla, 1972; Huys et al., 2011), laying the foundation for modeling behavior by Reinforcement Learning (RL; Sutton and Barto, 2018). To describe reward learning behaviors emerging from a reward learning task, Huys et al. (2013) introduced a prediction error RL approach to characterize decision making based on key behavioral phenotypes, including reward sensitivity and learning rate. In a PRT study, they identified reward sensitivity (instead of learning rate) as an important behavioral marker contributing to the abnormality in reward learning for MDD, with a lower reward sensitivity in MDD patients. Guo et al. (2023) extends the approach of Huys et al. (2013) to a more general semiparametric RL framework by replacing the scalar value of reward sensitivity with a nonlinear and nondecreasing reward sensitivity function. They revealed that the reward sensitivity for the MDD patients is not uniformly lower than the healthy subjects, but only when subjects receive sufficient rewards.
Furthermore, Guo et al. (2023) noted a floor and ceiling effect in the relationship between choices and expected rewards, attributed to the non-linearity of the reward sensitivity function, indicating that the decision-making process may not adhere to a single RL model. Several other studies also support that subjects employ multiple learning strategies in decision-making, as highlighted by Worthy et al. (2013); Iigaya et al. (2018); Ashwood et al. (2022). In particular, Ashwood et al. (2022) demonstrates that the perceptual decision-making process switches between multiple interleaved strategies. They suggest that the decision process involves a mixture of ‘engaged’ and ‘lapse’ strategies. During the ‘engaged’ strategy, a subject makes choices following a generalized linear model with softmax. Conversely, during the ‘lapse’ strategy, the subject ignores the stimulus and makes choices based on a fixed probability. The dynamics of strategy-switching are modeled by a hidden Markov model (HMM). Alternatively, Chen et al. (2021) considers ‘engaged’ as exploitation and ‘lapse’ as exploration, aligning with the exploitation-exploration trade-off in RL (Sutton and Barto, 2018). During the exploration stages, each choice is assigned equal probabilities.
@@
In this work, we propose an RL-HMM framework to characterize reward-based decision-making with alternating learning strategies between two distinct decision-making processes: in the ‘engaged’ state, subjects make decisions based on the RL model, while in the ‘lapse’ state, subjects make decisions through random choices with equal probabilities. Our approach generalizes the RL model introduced in Huys et al. (2013) and Guo et al. (2023) from finite discrete RL states to infinite but bounded state space. We employ HMM with time-varying transition probabilities to model the learning strategy switching, extending the HMM framework in Ashwood et al. (2022), where the transition probabilities are assumed to be time-invariant. In our approach, the time-varying transition probabilities are modeled nonparametrically with fused lasso regularization (Tibshirani et al., 2005) or trend filtering (Tibshirani, 2014). The proposed approach is applied to the Probabilistic Reward Task (Pizzagalli et al., 2005) in the EMBARC study (Trivedi et al., 2016), a large-scale randomized clinical trial recruiting participants diagnosed with MDD and healthy controls (details in Section 4.1).
@@ 2 Methods
Consider a group of $n$ subjects. The dataset comprises decision-making processes of each subject, (i.e., $\dots,S_{it}\rightarrow A_{it}\rightarrow R_{it}\rightarrow S_{it+1}\rightarrow A_{it+1}\rightarrow R_{it+1},\dots$), where $S_{it}$, $A_{it}$, and $R_{it}$ represent the random variables of state, action, and reward for the $i$-th subject at time $t$, respectively. Throughout this paper, we assume that the states reside within a bounded state space denoted as $\mathcal{S}$. We consider a discrete action space $\mathcal{A}$ with $D$ available options (i.e. $\mathcal{A}=\{0,1,\dots,D-1\}$). The reward $R_{it}$ is generated based on a function involving $S_{it}$, $A_{it}$, and $t$. We assume that $R_{it}$ is bounded and can be discrete or continuous. In the task of modeling human reward learning abilities in behavioral experiments, rewards are provided without delay.
@@ 2.1 Decision-making with reinforcement learning
Reinforcement Learning (RL) is a computational framework that can be used to model subjects’ responses to rewards in behavioral tasks. In this framework, subjects learn the underlying reward generating mechanism and subsequently make decisions based on their acquired knowledge.
Define $Q_{i}^{t}(a,s)$ as the expected reward for taking action $a$ at state $s$ for the $i$th participant at the $t$th trial. Assume that the expected reward can be approximated by $\widetilde{Q}_{i}^{t}(a,s)=\bm{\phi}(a,s)^{\top}\bm{\delta}_{i}^{t}$, where $\bm{\phi}(a,s)$ represents the basis vector associated with the state $s\in\mathcal{S}$ for action $a\in\mathcal{A}$. After observing the reward $R_{it}=r_{it}$ at trial $t$, the participant updates the expected reward $\widetilde{Q}_{i}^{t}(a,s)$ by adjusting it in the direction of $r_{it}-\widetilde{Q}_{i}^{t}(a,s)$. The corresponding update of the coefficients follows:
@@
where $\beta$ is the learning rate. Note that (1) is also known as the Rescorla-Wagner equation (Rescorla, 1972). It can be viewed as a one-step stochastic gradient descent, where we minimize the risk associated with the following mean squared error:
@@
$Q_{i}^{t+1}(a_{it},s_{it})=\beta_{it}r_{it}+(1-\beta_{it})Q_{i}^{t}(a_{it},s_{it}),$ (3)
where $\beta_{it}=\beta\bm{\phi}(a_{it},s_{it})^{\top}\bm{\phi}(a_{it},s_{it})$. To ensure that the expected reward at $t+1$ is the weighted sum of the obtained reward and the expected reward at $t$, the learning rate should satisfy $10<\beta_{it}<$ 1. As a note, if $s$ is discrete and we choose $\phi(a,s)$ as indicator functions on all discrete values in $\mathcal{A}\times\mathcal{S}$, then $\bm{\phi}(a_{it},s_{it})^{\top}\bm{\phi}(a_{it},s_{it})\equiv 1$ and $\beta_{it}\equiv\beta$ in (3). Alternatively, if $s$ is continuous, we choose $\phi(a,s)=(I(a=0),I(a=1))^{\top}\otimes\psi(s)$, where $\psi(s)$ can be a linear basis or B-splines. We further assume that $\bm{\alpha}:=\bm{\delta}^{1}=\bm{\delta}_{i}^{1}$. Define $\mathcal{H}_{it}=\{S_{ij},A_{ij},R_{ij}\}_{j=1}^{t-1}$ as the observed history for the $i$-th subject up to time $t$. We assume that the action $A_{it}$ follows the model
@@
where $\rho>0$ is the reward sensitivity. A higher reward sensitivity indicates that subjects’ actions will depend more on the expected rewards they learned. Here, $\nu_{a}$ $(a=0,\dots,D-1)$ are time-invariant intercepts. To ensure identifiability, we let $\nu_{0}\equiv 0$, and $\nu_{1}=\dots=\nu_{D-1}=0$ implies that the subject is making decisions entirely based on the expected reward.
2.2 Decision-making with state switching learning strategies
Several studies provide evidence that subjects employ multiple learning strategies for decision-making (Worthy et al., 2013; Iigaya et al., 2018; Ashwood et al., 2022). In this work, we consider that subjects exhibit two distinct learning strategies: the ‘engaged’ state (represented as $U_{it}=1$), in which subjects make decisions based on the RL model (Huys et al., 2013; Guo et al., 2023), and the ‘lapse’ state (represented as $U_{it}=0$), where subjects make decisions through random choices, with an equal probability of $1/D$ for selecting one of the $D$ available actions (Chen et al., 2021). Here, $U_{it}\in\{0,1\}$ represents a latent variable that indicates the learning strategy used by the subject $i$ during trial $t$. We introduce a novel RL-HMM framework for the analysis of perceptual decision-making behavior, employing a combination of hidden Markov model (HMM) and reinforcement learning. The proposed RL-HMM model is:

### v2

# HMM for Discovering Decision-Making Dynamics Using Reinforcement Learning Experiments

@@ Abstract
Major depressive disorder (MDD), a leading cause of years of life lived with disability, presents challenges in diagnosis and treatment due to its complex and heterogeneous nature. Emerging evidence indicates that reward processing abnormalities may serve as a behavioral marker for MDD. To measure reward processing, patients perform computer-based behavioral tasks that involve making choices or responding to stimulants that are associated with different outcomes, such as gains or losses in the laboratory. Reinforcement learning (RL) models are fitted to extract parameters that measure various aspects of reward processing (e.g., reward sensitivity) to characterize how patients make decisions in behavioral tasks. Recent findings suggest the inadequacy of characterizing reward learning solely based on a single RL model; instead, there may be a switching of decision-making processes between multiple strategies. An important scientific question is how the dynamics of strategies in decision-making affect the reward learning ability of individuals with MDD. Motivated by the probabilistic reward task (PRT) within the EMBARC study, we propose a novel RL-HMM framework for analyzing reward-based decision-making. Our model accommodates decision-making strategy switching between two distinct approaches under a hidden Markov model (HMM): subjects making decisions based on the RL model or opting for random choices. We account for continuous RL state space and allow time-varying transition probabilities in the HMM. We introduce a computationally efficient EM algorithm for parameter estimation and employ a nonparametric bootstrap for inference. Extensive simulation studies validate the finite-sample performance of our method. We apply our approach to the EMBARC study to show that MDD patients are less engaged in RL compared to the healthy controls, and engagement is associated with brain activities in the negative affect circuitry during an emotional conflict task. reinforcement learning; reward tasks; state-switching; behavioral phenotyping; mental health; brain-behavior association
@@ 1 Introduction
Despite extensive research and clinical efforts dedicated to developing pharmacological and behavioral treatments for mental disorders over the years, a significant setback is symptom-based diagnoses that assume discrete disease categories, which may contribute to low remission rates and inadequate treatment responses as it does not fully account for the heterogeneity and complexity of mental disorders (Rush et al.,, 2006). In response to the limitations inherent in traditional symptom-based diagnoses and to address significant between-patient heterogeneity, the National Institute of Mental Health (NIMH) advocates for a paradigm shift through the Research Domain Criteria (RDoC) initiative (Insel et al.,, 2010). This shift involves redefining mental disorders by identifying latent constructs derived from biological and behavioral measures across diverse domains of functioning (e.g., positive/negative affect, decision-making, social processing) at different levels (e.g., cells, genes, and brain circuits), instead of solely relying on the clinical symptoms. Behavioral data collected from tasks such as the probabilistic reward task (PRT; Pizzagalli et al.,, 2005) and the emotion conflict task (Etkin et al.,, 2006) is a highly valued resource within the framework of RDoC. Utilizing carefully designed behavioral tasks with appropriate data- or theory-driven computational methods (Huys et al.,, 2016), researchers can extract valuable behavioral markers that serve as indicators of mental health conditions. Moreover, there is a critical need to examine the associations between the brain and behavior, as demonstrated by Fonzo et al., (2019). This exploration facilitates understanding the underlying brain mechanisms during behavioral tasks, allowing for the identification of brain abnormalities related to mental disorders, and ultimately contributes to refining diagnostic frameworks and developing more effective therapeutic interventions.
@@
Reward learning and decision-making are pivotal in the etiology of mental function (Kendler et al.,, 1999). Reward processing refers to how individuals perceive, value, and pursue rewards (e.g., money, social approval). MDD patients may show abnormalities in reward processing, such as reduced motivation, anhedonia (loss of pleasure), or reduced sensitivity to positive rewards. To measure reward processing, patients perform computer-based behavioral tasks that involve making choices or responding to stimulants that are associated with different outcomes, such as gains or losses in the laboratory. We focus on characterizing human decision-making behaviors through reward tasks (e.g., PRT), wherein individuals make decisions by interacting with a reward-based system and continually adapting their decision-making mechanisms based on obtained rewards.
@@
Biological evidence suggests that fluctuations in human dopaminergic neurons indicate errors in predicting future rewarding events (Schultz et al.,, 1997). This supports the use of simple prediction error learning (comparing real and expected reward) to update the value functions of different choices (Rescorla,, 1972; Huys et al.,, 2011), laying the foundation for modeling behavior by Reinforcement Learning (RL; Sutton and Barto,, 2018). To describe reward learning behaviors emerging from a reward learning task, Huys et al., (2013) introduced a prediction error RL approach to characterize decision making based on key behavioral phenotypes, including reward sensitivity and learning rate. In a PRT study, they identified reward sensitivity (instead of learning rate) as an important behavioral marker contributing to the abnormality in reward learning for MDD, with a lower reward sensitivity in MDD patients. Guo et al., (2024) extends the approach of Huys et al., (2013) to a more general semiparametric RL framework by replacing the scalar value of reward sensitivity with a nonlinear and nondecreasing reward sensitivity function. They revealed that the reward sensitivity for the MDD patients is not uniformly lower than the healthy subjects, but only when subjects receive sufficient rewards. Furthermore, Guo et al., (2024) noted a floor and ceiling effect in the relationship between choices and expected rewards, attributed to the non-linearity of the reward sensitivity function, indicating that the decision-making process may not adhere to a single RL model. Several other studies also support that subjects employ multiple strategies in decision-making, as highlighted by Worthy et al., (2013); Iigaya et al., (2018); Ashwood et al., (2022). In particular, Ashwood et al., (2022) demonstrates that the perceptual decision-making process switches between multiple interleaved strategies. They suggest that the decision process involves a mixture of ‘engaged’ and ‘lapse’ strategies. During the ‘engaged’ strategy, a subject makes choices following a generalized linear model with softmax. Conversely, during the ‘lapse’ strategy, the subject ignores the stimulus and makes choices based on a fixed probability. The dynamics of strategy-switching are modeled by a hidden Markov model (HMM). Alternatively, Chen et al., (2021) considers ‘engaged’ as exploitation and ‘lapse’ as exploration, aligning with the exploitation-exploration trade-off in RL (Sutton and Barto,, 2018). During the exploration stages, each choice is assigned equal probabilities.
@@
In this work, we propose an RL-HMM framework to characterize reward-based decision-making with alternating strategies between two distinct decision-making processes: in the ‘engaged’ state, subjects make decisions based on the RL model, while in the ‘lapse’ state, subjects make decisions through random choices with equal probabilities. Our approach generalizes the RL model introduced in Huys et al., (2013) and Guo et al., (2024) from finite discrete RL states to infinite but bounded state space. We employ HMM with time-varying transition probabilities to model the decision-making strategy switching, extending the HMM framework in Ashwood et al., (2022), where the transition probabilities are assumed to be time-invariant. In our approach, the time-varying transition probabilities are modeled nonparametrically with fused lasso regularization (Tibshirani et al.,, 2005) or trend filtering (Tibshirani,, 2014). The proposed approach is applied to the Probabilistic Reward Task (Pizzagalli et al.,, 2005) in the EMBARC study (Trivedi et al.,, 2016), a large-scale randomized clinical trial recruiting participants diagnosed with MDD and healthy controls (details in Section 4.1). To the best of our knowledge, this study is the first to consider the decision-making dynamics between RL and other decision-making strategies in probabilistic reward tasks. Compared to Huys et al., (2013) and Guo et al., (2024), this work allows us to capture additional important features regarding the engagement patterns of both the MDD group and each individual subject in the task. We demonstrate the importance of parameters obtained from our model, such as the group engagement rates and individual engagement scores (defined in Section 2.2). The group engagement rates reveal different decision-making strategy patterns between MDD patients and healthy subjects. The individual engagement scores serve as crucial behavioral markers that may be associated with specific brain regions or networks. Understanding engagement patterns provides valuable insights into how patients interact with and adhere to treatment or intervention protocols in clinical trials. It can help healthcare providers and researchers optimize strategies to enhance patient involvement and ultimately improve treatment outcomes. Our proposed work aligns with the framework of imitation learning (Ross and Bagnell,, 2010) or behavioral cloning (Torabi et al.,, 2018), wherein a policy is learned by mimicking the behavior demonstrated by an expert from their observed decision-making process. One difference is that we do not assume participants follow an optimal policy that maximizes their rewards. A related line of research is inverse reinforcement learning (Abbeel and Ng,, 2004).
@@ 2 Methods
Consider a group of $n$ subjects. The dataset comprises decision-making processes of each subject, (i.e., $\dots,S_{it},\ A_{it},\ R_{it},\ S_{it+1},\ A_{it+1},\ R_{it+1},\dots$), where $S_{it}$, $A_{it}$, and $R_{it}$ represent the random variables of state, action, and reward for the $i$-th subject at time $t$, respectively. Define $\mathcal{H}_{it}=\{S_{ij},A_{ij},R_{ij}\}_{j=1}^{t-1}$ as the observed history for the $i$-th subject up to time $t$. Throughout this paper, we assume that the state $S_{it}$ is generated from an arbitrary non-trivial distribution of the transitions, denoted as $P(S_{it}\mid\mathcal{H}_{it})$, with these states confined within a bounded state space denoted as $\mathcal{S}$. One special case is the contextual bandit problem, where states are generated independently, i.e., $P(S_{it}\mid\mathcal{H}_{it})=P(S_{it})$. We consider a discrete action space $\mathcal{A}$ with $D$ available options (i.e. $\mathcal{A}=\{0,1,\dots,D-1\}$). The decision-making models are introduced in detail in Sections 2.1 to 2.2. The reward $R_{it}$ is generated based on a function involving $S_{it}$, $A_{it}$, and $t$ (i.e., the reward-generating distribution is $P_{t}(R_{it}\mid S_{it},A_{it})$), with the reward space $\mathcal{R}\subset[0,R_{\max}]$ being either discrete or continuous. The mechanisms that generate rewards can be either stochastic or deterministic. In this paper, our aim is to model human reward-based decision-making behavior in behavioral experiments where immediate rewards are provided at the end of each trial. Therefore, we do not consider downstream effects of rewards.
@@ 2.1 Decision-making with reinforcement learning
RL is a computational framework that can be used to model subjects’ responses to rewards in behavioral tasks. In this framework, subjects learn the underlying reward generating mechanism and subsequently make decisions based on their acquired knowledge.
Define $Q_{i}^{t}(a,s)$ as the expected reward for taking action $a$ at state $s$ for the $i$th participant at the $t$th trial. Assume that the expected reward can be approximated by $\widetilde{Q}_{i}^{t}(a,s)=\bm{\phi}(a,s)^{\top}\bm{\delta}_{i}^{t}$, where $\bm{\phi}(a,s)$ represents the basis vector associated with the state $s\in\mathcal{S}$ for action $a\in\mathcal{A}$, and $\bm{\delta}_{i}^{t}$ is the corresponding coefficient vector. After observing the reward $R_{it}=r_{it}$ at trial $t$, the participant updates the expected reward $\widetilde{Q}_{i}^{t}(a,s)$ by adjusting it in the direction of $r_{it}-\widetilde{Q}_{i}^{t}(a,s)$. The corresponding update of the coefficients follows:
@@
where $\beta$ is the learning rate. Note that (1) is also known as the Rescorla-Wagner equation (Rescorla,, 1972). It can be viewed as a one-step stochastic gradient descent, where we minimize the risk associated with the following mean squared error:
@@
$\widetilde{Q}_{i}^{t+1}(a_{it},s_{it})=\beta_{it}r_{it}+(1-\beta_{it})\widetilde{Q}_{i}^{t}(a_{it},s_{it}),$ (3)
where $\beta_{it}=\beta\bm{\phi}(a_{it},s_{it})^{\top}\bm{\phi}(a_{it},s_{it})$. To ensure that the expected reward at $t+1$ is the weighted sum of the obtained reward and the expected reward at $t$, the learning rate should satisfy $10<\beta_{it}<$ 1. As a note, if $s$ is discrete and we choose $\phi(a,s)$ as indicator functions on all discrete values in $\mathcal{A}\times\mathcal{S}$, then $\bm{\phi}(a_{it},s_{it})^{\top}\bm{\phi}(a_{it},s_{it})\equiv 1$ and $\beta_{it}\equiv\beta$ in (3). Alternatively, if $s$ is continuous, we choose $\phi(a,s)=(I(a=0),I(a=1))^{\top}\otimes\psi(s)$, where $\psi(s)$ can be a linear basis or B-splines. We further assume that $\bm{\alpha}:=\bm{\delta}^{1}=\bm{\delta}_{i}^{1}$. We assume that the action $A_{it}$ follows the model
@@
where $\rho>0$ is the reward sensitivity. A higher reward sensitivity indicates that subjects’ actions will depend more on the expected rewards they learned. Here, $\nu_{a}$ $(a=0,\dots,D-1)$ are time-invariant intercepts. To ensure identifiability, we let $\nu_{0}\equiv 0$, and $\nu_{1}=\dots=\nu_{D-1}=0$ implies that the subject is making decisions entirely based on the expected reward. The softmax decision-making model (4) is widely used in policy gradient methods (Sutton and Barto,, 2018) in RL because it is differentiable with respect to its parameters and its stochastic nature ensures exploration.
2.2 Decision-making with state switching strategies
Several studies provide evidence that subjects employ multiple strategies for decision-making (Worthy et al.,, 2013; Iigaya et al.,, 2018; Ashwood et al.,, 2022). In this work, we consider that subjects exhibit two distinct decision-making strategies: the ‘engaged’ state (represented as $U_{it}=1$), in which subjects make decisions based on the RL model (Huys et al.,, 2013; Guo et al.,, 2024), and the ‘lapse’ state (represented as $U_{it}=0$), where subjects make decisions through random choices, with an equal probability of $1/D$ for selecting one of the $D$ available actions (Chen et al.,, 2021). Here, $U_{it}\in\{0,1\}$ represents a latent variable that indicates the decision-making strategy used by the subject $i$ during trial $t$. We assume that in the behavioral decision-making framework, how subjects make decisions is independent of how they learn from observed (action, reward) pairs. Even in the ‘lapse’ phase, subjects still internally process reward updates by referencing the expected reward output, akin to having a computational mechanism in their brain. Engagement, however, depends on various factors such as mood, emotions, and potentially depression, which can vary significantly. For instance, a significant issue in mobile health is that some patients disengage from their studies due to fatigue induced by prompts. We introduce a novel RL-HMM framework for the analysis of perceptual decision-making behavior, employing a combination of hidden Markov model (HMM) and reinforcement learning. The proposed RL-HMM model is: